OpenStudy (anonymous):

whats next 3 terms for 12,14,16 and 224,112,56 and 3,6,9,12 also can u tell me how you did this :D

OpenStudy (anonymous):

well you can solve the first one by adding 2 to each term the second you are dividing each term by 2 the third one you are adding 3 to each term solve by using this information

OpenStudy (anonymous):

i don't understand adding each term ?? i am not sure wat term means

OpenStudy (anonymous):

a term is the number and for the first set 12+2=14+2=16 that is what I ment